# Robust Tail Index Estimation under Random Censoring via Minimum Density Power Divergence

### v5

@@ 6. Conclusion
In this work, we addressed the problem of robust estimation of the extreme value index under random right censoring, a setting frequently encountered in survival analysis, reliability engineering, and insurance. Although various estimators exist in the literature, including Hill-type methods and those based on Kaplan–Meier or Nelson–Aalen estimators, many remain sensitive to outliers or rely on strong parametric assumptions.
To overcome these limitations, we introduced a new estimator based on the Minimum Density Power Divergence (MDPD), a criterion well known for its robustness in the uncensored context. We carefully extended this methodology to handle randomly right-censored data, accounting for the inherent incompleteness in such settings. To the best of our knowledge, this represents the first application of the MDPD framework to tail index estimation under censoring, thereby filling a notable gap in the robust extreme value literature.
We established the consistency and asymptotic normality of the proposed estimator under standard regular variation assumptions. A comprehensive simulation study demonstrated its strong finite-sample performance, showing an advantageous balance between robustness and efficiency, particularly under model misspecification or contamination.
The practical relevance of the approach is illustrated using the insurance loss dataset, which satisfies the weak censoring condition ($p>1/2$). Here, the estimator performs in line with theoretical expectations and demonstrates clear robustness gains when contamination affects the latent losses before censoring. For the Australian AIDS survival dataset, which exhibits strong censoring ($p<1/2$), most tail index estimators fail to provide reliable results, highlighting the method’s limitations under such extreme conditions. The analysis also confirms that contamination introduced after censoring does not meaningfully perturb the latent distribution, and thus the MDPD framework should focus on pre-censoring contamination for coherent robustness assessment.
This contribution opens new perspectives for developing robust methods in censored extreme value models. It provides a foundation for future research, including adaptive threshold selection, models with dependent censoring, robust estimation under covariate-dependent censoring, and extensions to scenarios where the proportion of uncensored extremes is small ($p<1/2$).

### v6

@@ 6. Conclusion
In this work, we addressed the problem of robust estimation of the extreme value index under random right censoring, a framework that naturally arises in survival analysis, reliability engineering, and insurance applications. While a variety of estimators have been proposed in the literature—including Hill-type procedures and methods based on Kaplan–Meier or Nelson–Aalen estimators—many of them remain sensitive to outliers or rely on restrictive modeling assumptions, particularly in the presence of contamination.
To overcome these limitations, we introduced a new tail index estimator based on the Minimum Density Power Divergence (MDPD), a criterion well known for its robustness properties in uncensored settings. We carefully extended the MDPD methodology to the random right-censoring framework, explicitly accounting for the incomplete nature of the observed data. To the best of our knowledge, this work constitutes the first application of the MDPD framework to tail index estimation under random censoring, thereby filling an important gap in the robust extreme value literature.
Under standard regular variation assumptions and within the weak censoring regime, we established the consistency and asymptotic normality of the proposed estimator. A comprehensive simulation study confirmed its strong finite-sample performance, highlighting a favorable robustness–efficiency trade-off, especially under model misspecification and contamination.
The practical relevance of the approach was illustrated using an insurance loss dataset satisfying the weak censoring condition $(p>1/2).$ In this setting, the estimator behaves in accordance with the theoretical results and exhibits clear robustness gains when contamination affects the latent losses prior to censoring. In contrast, the analysis of the Australian AIDS survival dataset, characterized by strong censoring $(p\leq 1/2),$ underscores the intrinsic difficulties of tail inference in such regimes, where most existing estimators fail to deliver stable or reliable results. This empirical evidence also confirms that contamination acting after censoring does not correspond to a meaningful perturbation of the latent distribution, and therefore provides limited insight for robustness assessment in reconstruction-based approaches.
Overall, this contribution opens new perspectives for the development of robust inference tools in censored extreme value models. Promising directions for future research include adaptive threshold selection, extensions to dependent or covariate-dependent censoring mechanisms, and the design of robust estimators capable of handling severe censoring scenarios where the proportion of uncensored extremes is small $(p\leq 1/2).$
@@
7. Acknowledgements
The authors would like to thank the anonymous reviewers for their careful reading of the manuscript and their constructive comments, which helped improve the clarity and quality of this work. We also acknowledge helpful discussions with colleagues that contributed to the development of the theoretical and numerical aspects of the paper. Any remaining errors are the sole responsibility of the authors.